One-tailed test; variance is known, α = 0.05, n = 35 



1
Expert's answer
2021-02-24T07:09:42-0500

n = 35

α = 0.05

One-tailed test

For one-tailed test, the confidence coefficient is

(1 – α) = 1 -0.05 = 0.95

Because the critical region is spread on one side (right or left) of the distribution curve. One tail test has a single critical value.

From the t table, the critical t value for the degree of freedom 35, and the probability 0.95 is 1.6896.